what is log (x - 48) + logx =2?
Answer by nerdybill(5404)

Applying "log rules":
log (x - 48) + logx =2?
log (x-48)x = 2
(x-48)x = 10^2
x^2 - 48x = 100
x^2 - 48x - 100 = 0
(x-50)(x+2) = 0
x = {-2, 50}
Since log of negative numbers doesn't make sense, throw it out leaving:
x = 50